# Mutual Information Density of Massive MIMO Systems over Rayleigh-Product Channels

### 05 Jan 2024 v4

@@ C-B1 The evaluation of $U_{1}$
$|\frac{1}{M}\mathbb{E}\underline{a}{\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}}\underline{b}|\leq\frac{1}{M}\mathrm{Var}^{\frac{1}{2}}(a)\mathrm{Var}^{\frac{1}{2}}(b)={\mathcal{O}}(\frac{\sqrt{\operatorname{Tr}{\mathbb{A}}^{2}}}{M^{\frac{3}{2}}}),$ (124)
where $a=\operatorname{Tr}{\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}{\mathbb{H}}^{H}$ and $b=\operatorname{Tr}{\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}$. Considering the technique used in (100), we can evaluate the second term in (119) as
$-\frac{1}{M}\mathbb{E}\operatorname{Tr}{\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}\operatorname{Tr}{\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}{\mathbb{H}}^{H}\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}=-\frac{\mathbb{E}\operatorname{Tr}{\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}{\mathbb{H}}^{H}}{M}\mathbb{E}\underline{\operatorname{Tr}{\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}}\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}-\omega\mathbb{E}\operatorname{Tr}{\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}{\mathbb{H}}^{H}\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}+{\mathcal{O}}(\frac{\sqrt{\operatorname{Tr}{\mathbb{A}}^{2}}}{M^{\frac{3}{2}}}),$ (125)
where $\frac{\mathbb{E}\operatorname{Tr}{\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}{\mathbb{H}}^{H}}{M}\mathbb{E}\underline{\operatorname{Tr}{\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}}\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}={\mathcal{O}}(\frac{1}{M^{\frac{3}{2}}})$. By Poincarè-Nash inequality, we have
$|\mathbb{E}\frac{u^{2}\partial\nu_{n}^{{\mathbb{Y}},{\mathbb{Z}}}}{\partial Y_{j,i}}\frac{[{\mathbb{Z}}^{H}{\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}]_{j,i}}{M}|={\mathcal{O}}(\frac{u^{2}}{M}(\frac{\operatorname{Tr}{\mathbb{A}}^{2}}{M}+1)),~{}\mathrm{Var}^{\frac{1}{2}}(\frac{u^{2}\partial\nu_{n}^{{\mathbb{Y}},{\mathbb{Z}}}}{\partial Y_{j,i}}\frac{[{\mathbb{Z}}^{H}{\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}]_{j,i}}{M})={\mathcal{O}}(\frac{u^{2}(\frac{\operatorname{Tr}{\mathbb{A}}^{2}}{M^{\frac{1}{2}}}+1)}{M^{2}}),$ (126)
$|\mathbb{E}\frac{u^{3}\partial\beta_{n}^{{\mathbb{Y}},{\mathbb{Z}}}}{\partial Y_{j,i}}\frac{[{\mathbb{Z}}^{H}{\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}]_{j,i}}{M}|={\mathcal{O}}(\frac{u^{3}}{M^{2}}(\frac{\operatorname{Tr}{\mathbb{A}}^{2}}{M}+1)),~{}\mathrm{Var}^{\frac{1}{2}}(\frac{u^{3}\partial\beta_{n}^{{\mathbb{Y}},{\mathbb{Z}}}}{\partial Y_{j,i}}\frac{[{\mathbb{Z}}^{H}{\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}]_{j,i}}{M})={\mathcal{O}}(\frac{u^{3}(\frac{\operatorname{Tr}{\mathbb{A}}^{2}}{M^{\frac{1}{2}}}+1)}{M^{3}}),$
@@
$\mathrm{Var}(\frac{\operatorname{Tr}{\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}{\mathbb{H}}^{H}{\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}{\mathbb{H}}^{H}}{M})={\mathcal{O}}(\frac{\operatorname{Tr}{\mathbb{A}}^{4}}{M^{3}})={\mathcal{O}}(\frac{(\operatorname{Tr}{\mathbb{A}}^{2})^{2}}{M^{3}}),$ (127)
$\mathrm{Var}(\frac{\operatorname{Tr}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}^{2}{\mathbb{H}}^{H}{\mathbb{Q}}}{M})={\mathcal{O}}(\frac{\operatorname{Tr}{\mathbb{A}}^{4}}{M^{3}})={\mathcal{O}}(\frac{(\operatorname{Tr}{\mathbb{A}}^{2})^{2}}{M^{3}}).$
Therefore, by noticing $\operatorname{Tr}{\mathbb{A}}=0$, plugging (125) into (119), moving $-\omega\mathbb{E}\operatorname{Tr}{\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}{\mathbb{H}}^{H}\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}$ to the LHS of (119), and solving $\mathbb{E}\operatorname{Tr}{\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}{\mathbb{H}}^{H}\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}$, (119) can be rewritten as
$\mathbb{E}\operatorname{Tr}{\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}{\mathbb{H}}^{H}\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}$ $=-\frac{\jmath u\overline{\omega}n}{\sqrt{Mn}}\frac{\mathbb{E}\operatorname{Tr}{\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}{\mathbb{H}}^{H}{\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}{\mathbb{H}}^{H}}{M}\mathbb{E}\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}+\frac{\jmath u\overline{\omega}n}{\sqrt{Mn}}\frac{\mathbb{E}\operatorname{Tr}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}^{2}{\mathbb{H}}^{H}{\mathbb{Q}}}{M}\mathbb{E}\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}$ (128)
$=-\frac{\jmath u\overline{\omega}n}{\sqrt{Mn}}(X_{1}-X_{2})\mathbb{E}\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}+{\mathcal{O}}(\frac{\sqrt{\operatorname{Tr}{\mathbb{A}}^{2}}}{M^{\frac{3}{2}}}+\frac{u\operatorname{Tr}{\mathbb{A}}^{2}}{M^{\frac{3}{2}}}+\frac{u^{2}(1+\frac{\operatorname{Tr}{\mathbb{A}}^{2}}{M})}{M}+\frac{u^{3}(1+\frac{\operatorname{Tr}{\mathbb{A}}^{2}}{M})}{M^{2}}).$
@@
$X_{1}=\mathbb{E}\frac{1}{M}\sum_{i,j}Y^{*}_{j,i}[{\mathbb{Z}}^{H}{\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}{\mathbb{H}}^{H}{\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}]_{j,i}$ (129)
$=\mathbb{E}\{\frac{\operatorname{Tr}{\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}}{M}\frac{\operatorname{Tr}{\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}^{2}{\mathbb{H}}^{H}}{M}+\frac{\operatorname{Tr}{\mathbb{A}}}{M}\frac{\operatorname{Tr}{\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}{\mathbb{H}}^{H}{\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}}{M}$
$-\frac{\operatorname{Tr}{\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}}{M}\frac{\operatorname{Tr}{\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}{\mathbb{H}}^{H}{\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}{\mathbb{H}}^{H}}{M}-\frac{\operatorname{Tr}{\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}{\mathbb{H}}^{H}}{M}\frac{\operatorname{Tr}{\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}{\mathbb{H}}^{H}}{M}$
$-\frac{\operatorname{Tr}{\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}{\mathbb{H}}^{H}}{M}\frac{\operatorname{Tr}{\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}{\mathbb{H}}^{H}}{M}\}$
@@
$A_{1}=\frac{\omega\mathbb{E}\operatorname{Tr}{\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}^{2}{\mathbb{H}}^{H}}{M}+{\mathcal{O}}(\frac{\operatorname{Tr}{\mathbb{A}}^{2}}{M^{\frac{5}{2}}}).$ (130)
By noticing that $\operatorname{Tr}{\mathbb{A}}=0$, we can obtain that $A_{2}=0$. $A_{3}$ can be evaluated as
$A_{3}=-\omega X_{1}+{\mathcal{O}}(\frac{\operatorname{Tr}{\mathbb{A}}^{2}}{M^{\frac{5}{2}}}),$ (131)
@@
$A_{4}=-(\mathbb{E}a)^{2}+{\mathcal{O}}(\frac{\operatorname{Tr}{\mathbb{A}}^{2}}{M^{3}})={\mathcal{O}}(\frac{\operatorname{Tr}{\mathbb{A}}^{2}}{M^{3}}),$ (132)
where $a=\frac{\operatorname{Tr}{\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}{\mathbb{H}}^{H}}{M}$ and the last equality follows from $\mathbb{E}a={\mathcal{O}}(\frac{\operatorname{Tr}{\mathbb{A}}^{2}}{M^{5}})$, which can be derived by a similar approach as (122). Define $c=\frac{\operatorname{Tr}{\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}{\mathbb{H}}^{H}}{M}$ and $d=\frac{\operatorname{Tr}{\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}{\mathbb{H}}^{H}}{M}$. By (61) in Lemma 2, we have $\mathrm{Var}(c)={\mathcal{O}}(\frac{\operatorname{Tr}{\mathbb{A}}^{2}}{M^{3}})={\mathcal{O}}(\frac{1}{M})$ and $\mathrm{Var}(d)={\mathcal{O}}(\frac{\operatorname{Tr}{\mathbb{A}}^{2}}{M^{3}})={\mathcal{O}}(\frac{1}{M})$. Therefore, $A_{5}$ can be handled by
$A_{5}=-\mathbb{E}c\mathbb{E}d+\mathrm{Cov}(c,d)={\mathcal{O}}(\frac{1}{M^{3}})+{\mathcal{O}}(\frac{\operatorname{Tr}{\mathbb{A}}^{2}}{M^{3}}),$ (133)
where $\mathbb{E}c$ and $\mathbb{E}d$ can be evaluated by the same approach as (122). By moving $\omega X_{1}$ from $A_{3}$ to the LHS of (129) and using the evaluation of $A_{1}$ in (130), we can solve $X_{1}$ as
$X_{1}=\overline{\omega}A_{1}+{\mathcal{O}}(\frac{\operatorname{Tr}{\mathbb{A}}^{2}}{M^{\frac{5}{2}}})=\frac{\omega\overline{\omega}\mathbb{E}\operatorname{Tr}{\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}^{2}{\mathbb{H}}^{H}}{M}+{\mathcal{O}}(\frac{\operatorname{Tr}{\mathbb{A}}^{2}}{M^{\frac{5}{2}}}).$ (134)

### 09 Apr 2024 v5

@@ C-B1 The evaluation of $U_{1}$
$|\frac{1}{M}\mathbb{E}[\underline{a}{\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}}\underline{b}]|\leq\frac{1}{M}\mathrm{Var}^{\frac{1}{2}}(a)\mathrm{Var}^{\frac{1}{2}}(b)={\mathcal{O}}(\frac{\sqrt{\operatorname{Tr}({\mathbb{A}}^{2})}}{M^{\frac{3}{2}}}),$ (124)
where $a=\operatorname{Tr}({\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}{\mathbb{H}}^{H})$ and $b=\operatorname{Tr}({\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H})$. Considering the technique used in (100), we can evaluate the second term in (119) as
$-\frac{1}{M}\mathbb{E}[\operatorname{Tr}({\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H})\operatorname{Tr}({\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}{\mathbb{H}}^{H})\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}]=-\frac{\mathbb{E}[\operatorname{Tr}({\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}{\mathbb{H}}^{H})]}{M}\mathbb{E}[\underline{\operatorname{Tr}({\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H})}\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}]-\omega\mathbb{E}[\operatorname{Tr}({\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}{\mathbb{H}}^{H})\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}]+{\mathcal{O}}(\frac{\sqrt{\operatorname{Tr}({\mathbb{A}}^{2})}}{M^{\frac{3}{2}}}),$ (125)
where $\frac{\mathbb{E}[\operatorname{Tr}({\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}{\mathbb{H}}^{H})]}{M}\mathbb{E}[\underline{\operatorname{Tr}({\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H})}\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}]={\mathcal{O}}(\frac{1}{M^{\frac{3}{2}}})$. By Poincarè-Nash inequality, we have
$\Biggl{|}\mathbb{E}\Biggl{[}\frac{u^{2}\partial\nu_{n}^{{\mathbb{Y}},{\mathbb{Z}}}}{\partial Y_{j,i}}\frac{[{\mathbb{Z}}^{H}{\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}]_{j,i}}{M}\Biggr{]}\Biggr{|}={\mathcal{O}}\Bigl{(}\frac{u^{2}}{M}(\frac{\operatorname{Tr}({\mathbb{A}}^{2})}{M}+1)\Bigr{)},~{}\mathrm{Var}^{\frac{1}{2}}(\frac{u^{2}\partial\nu_{n}^{{\mathbb{Y}},{\mathbb{Z}}}}{\partial Y_{j,i}}\frac{[{\mathbb{Z}}^{H}{\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}]_{j,i}}{M})={\mathcal{O}}\Bigl{(}\frac{u^{2}(\frac{\operatorname{Tr}({\mathbb{A}}^{2})}{M^{\frac{1}{2}}}+1)}{M^{2}}\Bigr{)},$ (126)
$\Biggl{|}\mathbb{E}\Biggl{[}\frac{u^{3}\partial\beta_{n}^{{\mathbb{Y}},{\mathbb{Z}}}}{\partial Y_{j,i}}\frac{[{\mathbb{Z}}^{H}{\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}]_{j,i}}{M}\Biggr{]}\Biggr{|}={\mathcal{O}}\Bigl{(}\frac{u^{3}}{M^{2}}(\frac{\operatorname{Tr}({\mathbb{A}}^{2})}{M}+1)\Bigr{)},~{}\mathrm{Var}^{\frac{1}{2}}(\frac{u^{3}\partial\beta_{n}^{{\mathbb{Y}},{\mathbb{Z}}}}{\partial Y_{j,i}}\frac{[{\mathbb{Z}}^{H}{\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}]_{j,i}}{M})={\mathcal{O}}\Bigl{(}\frac{u^{3}(\frac{\operatorname{Tr}({\mathbb{A}}^{2})}{M^{\frac{1}{2}}}+1)}{M^{3}}\Bigr{)},$
@@
$\mathrm{Var}\Bigl{(}\frac{\operatorname{Tr}{\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}{\mathbb{H}}^{H}{\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}{\mathbb{H}}^{H}}{M}\Bigr{)}={\mathcal{O}}\Bigl{(}\frac{\operatorname{Tr}({\mathbb{A}}^{4})}{M^{3}}\Bigr{)}={\mathcal{O}}\Bigl{(}\frac{(\operatorname{Tr}({\mathbb{A}}^{2}))^{2}}{M^{3}}\Bigr{)},$ (127)
$\mathrm{Var}\Bigl{(}\frac{\operatorname{Tr}({\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}^{2}{\mathbb{H}}^{H}{\mathbb{Q}})}{M}\Bigr{)}={\mathcal{O}}\Bigl{(}\frac{\operatorname{Tr}({\mathbb{A}}^{4})}{M^{3}}\Bigr{)}={\mathcal{O}}\Bigl{(}\frac{(\operatorname{Tr}({\mathbb{A}}^{2}))^{2}}{M^{3}}\Bigr{)}.$
Therefore, by noticing $\operatorname{Tr}({\mathbb{A}})=0$, plugging (125) into (119), moving $-\omega\mathbb{E}[\operatorname{Tr}({\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}{\mathbb{H}}^{H})\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}]$ to the LHS of (119), and solving $\mathbb{E}[\operatorname{Tr}({\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}{\mathbb{H}}^{H})\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}]$, (119) can be rewritten as
$\mathbb{E}[\operatorname{Tr}({\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}{\mathbb{H}}^{H}\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}})]$ $=-\frac{\jmath u\overline{\omega}n}{\sqrt{Mn}}\frac{\mathbb{E}[\operatorname{Tr}({\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}{\mathbb{H}}^{H}{\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}{\mathbb{H}}^{H})]}{M}\mathbb{E}[\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}]+\frac{\jmath u\overline{\omega}n}{\sqrt{Mn}}\frac{\mathbb{E}[\operatorname{Tr}({\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}^{2}{\mathbb{H}}^{H}{\mathbb{Q}})]}{M}\mathbb{E}[\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}]$ (128)
$=-\frac{\jmath u\overline{\omega}n}{\sqrt{Mn}}(X_{1}-X_{2})\mathbb{E}[\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}]+{\mathcal{O}}\Bigl{(}\frac{\sqrt{\operatorname{Tr}({\mathbb{A}}^{2})}}{M^{\frac{3}{2}}}+\frac{u\operatorname{Tr}({\mathbb{A}}^{2})}{M^{\frac{3}{2}}}+\frac{u^{2}(1+\frac{\operatorname{Tr}({\mathbb{A}}^{2})}{M})}{M}+\frac{u^{3}(1+\frac{\operatorname{Tr}({\mathbb{A}}^{2})}{M})}{M^{2}}\Bigr{)}.$
@@
$X_{1}=\frac{1}{M}\sum_{i,j}\mathbb{E}[Y^{*}_{j,i}[{\mathbb{Z}}^{H}{\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}{\mathbb{H}}^{H}{\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}]_{j,i}]$ (129)
$=\mathbb{E}\Bigl{[}\frac{\operatorname{Tr}({\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H})}{M}\frac{\operatorname{Tr}({\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}^{2}{\mathbb{H}}^{H})}{M}+\frac{\operatorname{Tr}({\mathbb{A}})}{M}\frac{\operatorname{Tr}({\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}{\mathbb{H}}^{H}{\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H})}{M}$
$-\frac{\operatorname{Tr}({\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H})}{M}\frac{\operatorname{Tr}({\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}{\mathbb{H}}^{H}{\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}{\mathbb{H}}^{H})}{M}-\frac{\operatorname{Tr}({\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}{\mathbb{H}}^{H})}{M}\frac{\operatorname{Tr}({\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}{\mathbb{H}}^{H})}{M}$
$-\frac{\operatorname{Tr}({\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}{\mathbb{H}}^{H})}{M}\frac{\operatorname{Tr}({\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}{\mathbb{H}}^{H})}{M}\Bigr{]}$
@@
$A_{1}=\frac{\omega\mathbb{E}[\operatorname{Tr}({\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}^{2}{\mathbb{H}}^{H})]}{M}+{\mathcal{O}}\Bigl{(}\frac{\operatorname{Tr}({\mathbb{A}}^{2})}{M^{\frac{5}{2}}}\Bigr{)}.$ (130)
By noticing that $\operatorname{Tr}({\mathbb{A}})=0$, we can obtain that $A_{2}=0$. $A_{3}$ can be evaluated as
$A_{3}=-\omega X_{1}+{\mathcal{O}}\Bigl{(}\frac{\operatorname{Tr}({\mathbb{A}}^{2})}{M^{\frac{5}{2}}}\Bigr{)},$ (131)
@@
$A_{4}=-(\mathbb{E}[a])^{2}+{\mathcal{O}}(\frac{\operatorname{Tr}({\mathbb{A}}^{2})}{M^{3}})={\mathcal{O}}\Bigl{(}\frac{\operatorname{Tr}({\mathbb{A}}^{2})}{M^{3}}\Bigr{)},$ (132)
where $a=\frac{\operatorname{Tr}({\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}{\mathbb{H}}^{H})}{M}$ and the last equality follows from $\mathbb{E}[a]={\mathcal{O}}(\frac{\operatorname{Tr}({\mathbb{A}}^{2})}{M^{5}})$, which can be derived by a similar approach as (122). Define $c=\frac{\operatorname{Tr}({\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}{\mathbb{H}}^{H})}{M}$ and $d=\frac{\operatorname{Tr}({\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}{\mathbb{H}}^{H})}{M}$. By (61) in Lemma 2, we have $\mathrm{Var}(c)={\mathcal{O}}(\frac{\operatorname{Tr}({\mathbb{A}}^{2})}{M^{3}})={\mathcal{O}}(\frac{1}{M})$ and $\mathrm{Var}(d)={\mathcal{O}}(\frac{\operatorname{Tr}({\mathbb{A}}^{2})}{M^{3}})={\mathcal{O}}(\frac{1}{M})$. Therefore, $A_{5}$ can be handled by
$A_{5}=-\mathbb{E}[c]\mathbb{E}[d]+\mathrm{Cov}(c,d)={\mathcal{O}}(\frac{1}{M^{3}})+{\mathcal{O}}\Bigl{(}\frac{\operatorname{Tr}({\mathbb{A}}^{2})}{M^{3}}\Bigr{)},$ (133)
where $\mathbb{E}[c]$ and $\mathbb{E}[d]$ can be evaluated by the same approach as (122). By moving $\omega X_{1}$ from $A_{3}$ to the LHS of (129) and using the evaluation of $A_{1}$ in (130), we can solve $X_{1}$ as
$X_{1}=\overline{\omega}A_{1}+{\mathcal{O}}\Bigl{(}\frac{\operatorname{Tr}({\mathbb{A}}^{2})}{M^{\frac{5}{2}}}\Bigr{)}=\frac{\omega\overline{\omega}\mathbb{E}\operatorname{Tr}({\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}^{2}{\mathbb{H}}^{H})}{M}+{\mathcal{O}}\Bigl{(}\frac{\operatorname{Tr}({\mathbb{A}}^{2})}{M^{\frac{5}{2}}}\Bigr{)}.$ (134)